OpenStudy (anonymous):

Find the area of the triangle that has a base of 5 in. and a height of 3 3/4 in.

OpenStudy (jdoe0001):

area of a triangle = 1/2 x base x height

OpenStudy (anonymous):

@jdoe0001 i need help got 8 and 3/4 is that correct

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2} \times 5 \times 3\cfrac{3}{4}\\ \color{blue}{3\cfrac{3}{4} \implies \cfrac{15}{4} \text{thus then}}\\ \cfrac{5}{2}\times \cfrac{15}{4}\)

OpenStudy (anonymous):

\[5 X4=20\]

OpenStudy (anonymous):

\[15 x 2= 30\]

OpenStudy (anonymous):

\[20/30=2/3\]

OpenStudy (anonymous):

this was not one of the answer choices

OpenStudy (jdoe0001):

hmmm you seem to be dividing fractions instead, in a fraction multiplication, you multiply straight across \(\bf \cfrac{5 }{2}\frac{\Longleftrightarrow \times}{\Longleftrightarrow \times} \cfrac{15}{4}\)

OpenStudy (jdoe0001):

\(\cfrac{5}{2}\times \cfrac{15}{4} \implies \cfrac{75}{8}\)

OpenStudy (jdoe0001):

which you can also write as a mixed number of \(\bf 9\cfrac{3}{8}\)

OpenStudy (anonymous):

the answer choices is A)9 3/8 B)18 3/4 C)8 3/4 D)17 1/2

OpenStudy (anonymous):

o ok thank you so i had got it wrong at first thank you